OpenStudy (anonymous):

a tennis ball is served from a height of 2.4m from the ground at 30 m/s the net is 12m away and 0.9m high will the ball clear the net and where it will drop?

OpenStudy (imtiaz7):

make it s simple AS possible,in the horizontal direction,we do not have any acceleration so only horizontal speed=30m/s so speed=distance/time so time=0.4 seconds that is the time taken for ball to cover a horizontal range of 12m so next we have acceleration g in the vertical direction so but no initial veleocity to begin with as everythiong in horizontal later u=0 s=1/2at^2 so we have time 0.4 a=g=10 then displacement OR DISTNCE FROM THE GROUND LEVEL(HEIGHT) should be greater than the nets height(so that it is avbove it) that u got by calculation so now i would be happy if u said whether the ball crosses or not